Determine the pH of each of the following solutions.

Ka acetic acid is 1.8*10^-5.

Kb ammonia is 1.76*10^-5

Part A.

0.20 M NH4Cl

Part B.

0.19 M NaC2H3O2

Part C.

0.12 M NaCl

Thank you!
Answer #1

part A) NH4Cl is salt of weak base

pOH =1/2 [pKw + pKb + log C]

pKb of NH3 = - log Kb = - log 1.76 x 10-5 = 4.75

pOH = 1/2 [14 + 4.75 + log 0.20]

pOH = 9.02

pH = 14 - 9.02

pH = 4.98

part B) NaC2H3O2 is salt of weak acid

pH = 1/2 [pKw + pKa + log C]

pKa = - log Ka = - log 1.8 x 10-5 = 4.74

pH = 1/2 [14 + 4.74 + log 0.19]

pH = 9.01

part C) NaCl is salt of strong acid and strong base

for such salts pH always 7.0 only

pH = 7.0